# ArtistMus: A Globally Diverse, Artist-Centric Benchmark for Retrieval-Augmented Music Question Answering

### v1

@@ 1.  Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across diverse domains, yet their effectiveness in music-related tasks remains limited due to sparse music-specific coverage in pretraining data. This limitation is particularly evident in Music Question Answering (MQA), where users seek detailed information about artists, genres, instruments, and music history—knowledge that general-purpose LLMs often lack or hallucinate (ramoneda2024role).
Traditional approaches to domain adaptation rely on fine-tuning LLMs with domain-specific data (jeong2024fine; sahoo2024large). However, this faces significant challenges: securing high-quality training data, escalating computational costs, and the persistent difficulty of continuously updating model knowledge as new music emerges. Retrieval Augmented Generation (RAG) (lewis2021rag) offers a compelling alternative by grounding model outputs in external knowledge bases, enabling dynamic access to specialized information without retraining.
@@
Despite RAG’s success in domains such as medicine (miao2024integrating) and law (wiratunga2024cbr), its application in music remains underexplored due to the absence of music-specific retrieval resources. Existing knowledge bases like Wikipedia lack the density and structure needed for effective MQA. Similarly, current music benchmarks (weck2024muchomusic; yuan2024chatmusician; ramoneda2024role) focus primarily on music theory or multimodal understanding, leaving a critical gap in evaluating artist-centric knowledge—information central to everyday music discovery and consumption (lee2010analysis; doh2024music).
@@
Using these resources, we conduct systematic experiments showing that retrieval-augmented approaches markedly improve music QA performance. Open-source models equipped with MusWikiDB-based retrieval achieve substantial gains in factual accuracy, performing on par with or exceeding proprietary systems. Moreover, RAG-style training on (context, question, answer) triples further enhances factual recall and contextual reasoning. The framework generalizes well to out-of-domain benchmarks and yields more accurate and efficient retrieval than the general Wikipedia corpus. [^1] Code and data at https://anonymous.4open.science/r/MusWikiDB-and-ArtistMus
@@ 2.1.  Music Question Answering Benchmarks
Question Answering (QA) is a core task in Information Retrieval, involving the provision of accurate answers to given questions by retrieving relevant information from document collections (allam2012question). While open-domain QA systems operate over broad knowledge bases (rajpurkar2016squad; karpukhin2020dense), domain-specific QA targets specialized fields such as medicine (pal2022medmcqa), law (chalkidis2021lexglue), or music, where both documents and queries are confined to domain-specific knowledge.
Several recent benchmarks have emerged to evaluate LLM performance on music-related tasks. MuChoMusic (weck2024muchomusic) presents 1,187 audio-based multiple-choice questions assessing musical knowledge and reasoning. MusicTheoryBench (yuan2024chatmusician) contains 372 expert-validated questions evaluating advanced music theory knowledge. TrustMus (ramoneda2024role) comprises 400 questions across four domains—People, Instruments & Technology, Genres/Forms/Theory, and Culture & History—derived from The Grove Dictionary Online (sadie2001new). ZIQI-Eval (li2024music) offers 14,000 comprehension tasks spanning 10 major topics and 56 subtopics.
However, a critical gap exists in existing benchmarks: the inadequate representation of artist-centric metadata—information crucial for everyday music listening contexts (lee2010analysis; doh2024music). Current text-only QA benchmarks lack comprehensive coverage of details that listeners frequently seek, such as complete discographies, collaboration networks, creative evolution across albums, and career achievements. This disparity between current MQA capabilities and practical information demands motivates the development of artist-focused evaluation resources.
@@
Domain adaptation of LLMs traditionally relies on fine-tuning with domain-specific data (jeong2024fine; sahoo2024large; satterfield2024fine). However, this approach faces challenges: securing high-quality training data, escalating computational costs with model scale, and difficulties in continuously updating model knowledge.
Retrieval Augmented Generation (RAG) (lewis2021rag) offers an alternative by augmenting LLMs with external knowledge retrieval. Rather than relying solely on parametric memory, RAG retrieves relevant passages from external databases during inference to ground responses in factual context. Prior work demonstrates that RAG improves factual accuracy, reduces hallucination, enables transparent source verification, and provides greater scalability and cost-efficiency compared to fine-tuning alone (borgeaud2022improving; yasunaga2022deep; wang2023self).
While RAG has been successfully applied in general-purpose (ram2023context) and domain-specific settings such as law (wiratunga2024cbr) and medicine (miao2024integrating), its application in the music domain remains underexplored. Music QA requires both factual recall (e.g., artist names, compositions, release dates) and contextual reasoning (e.g., cultural influences, historical relations), making it an ideal testbed for evaluating RAG’s effectiveness in culturally rich, knowledge-intensive domains.
Recent work has also explored RAG-style fine-tuning, where models are trained on (context, question, answer) triples rather than traditional (question, answer) pairs (shuster2021retrieval). This approach explicitly teaches models to integrate supporting passages, potentially improving their ability to leverage external evidence. However, systematic evaluation of this strategy within specialized cultural domains like music has been limited.
@@ 3.  Dataset
Table 3: MusWikiDB and Wikipedia Corpus (karpukhin2020dense) statistics.
@@ 3.1.  MusWikiDB
To construct the RAG database from MusWikiDB_raw, the text was segmented into passages of up to 256 tokens with a 10% overlap between adjacent segments, based on the ablation study in Section 4.3, in order to preserve contextual continuity. For first-stage retrieval, we employed BM25 (robertson1994some), a classical yet highly effective algorithm for text relevance ranking, to build an efficient index for MusWikiDB. This enabled fast retrieval of relevant information during RAG-based inference, thereby improving both the accuracy and contextual relevance of the generated answers. The resulting MusWikiDB serves as a scalable and up-to-date knowledge base, enhancing RAG performance in MQA tasks and enabling the system to address complex, domain-specific music-related questions with greater precision and context.
Table 3 compares our proposed MusWikiDB with the Wikipedia corpus (karpukhin2020dense). While MusWikiDB contains fewer pages (0.14M vs 3.2M) and has a smaller vocabulary size (7.5M vs 21.5M), it consists exclusively of music-specialized text information.
Example Questions of ArtistMus Metadata: Artist: Thad Jones Country: United States Debut Year: 1954 Genre: Jazz Topic: Biography Factual Question: "Where was Thad Jones born?" A. New York City, New York B. Copenhagen, Denmark C. Des Moines, Iowa D. Pontiac, Michigan Answer: D Contextual Question: "What might have influenced Thad Jones’s decision to move to Copenhagen in 1979?" A. The chance to collaborate with Thelonious Monk on a new album B. The opportunity to lead The Danish Radio Big Band and transform it into a world-class ensemble C. The desire to escape the competitive music scene in New York City D. The need to focus solely on his teaching career at the Royal Danish Conservatory Answer: B
@@ 3.2.  ArtistMus
While previous benchmarks have mainly addressed general music understanding, none has focused on music metadata, particularly the artist, which is crucial in music listening contexts (lee2010analysis; doh2024music). Therefore, we introduce ArtistMus, a multiple-choice benchmark dataset on music artists, to evaluate the performance of LLMs in artist-related QA using artist-specific data from MusWikiDB.
@@ 3.2.1.  Data Collection
For genre normalization (schreiber2015improving), all genre labels were converted to lowercase, and spaces, hyphens (-), and slashes (/) were removed. We initially collected 48 root genres following (schreiber2016genre). After filtering to retain data corresponding to the 300 most frequent genres, each was mapped to one of 20 final genre labels.
To extract artists’ regional information, we provided the first paragraph (abstract) of each Wikipedia page to Llama 3.1 8B Instruct (grattafiori2024llama), prompting it to identify the country associated with each artist. The list of valid countries was obtained from the pycountry [^2] https://pypi.org/project/pycountry/ library.
@@ 3.2.2.  Data Selection
Figure 2 illustrates the regional distribution of the artists included in ArtistMus. Unlike many existing datasets that are heavily biased toward the United States and Europe, ArtistMus deliberately incorporates artists from 163 countries (or regions), ensuring broad global diversity and representation. Approximately 43% of the artists originate from outside the US and Europe, indicating balanced representation beyond Western and European cultural centers.
@@
For each artist, we generated one factual and one contextual question to evaluate the LLM’s factual accuracy and contextual reasoning ability. To construct these questions, we provided GPT-4o (achiam2023gpt) with the corresponding section text. Factual questions target verifiable details such as dates, names, and events, whereas contextual questions require reasoning or synthesis across multiple pieces of information within the passage.
@@ 3.2.3.  Data Validation
To ensure benchmark quality, we applied a two-stage validation process. We validate the generated questions based on two criteria: Music Relevance and Faithfulness. For Music Relevance, questions that did not pertain to musical aspects were excluded except important details such as the artist’s birthplace. For Faithfulness, LLM was asked to verify whether the question and answer could be derived from the provided text. Finally, 1,000 multiple-choice questions passing final human validation were generated. We adjusted the placement of correct answers to ensure an even distribution, assigning 250 instances to each option.
@@ 4.1.  Benchmarks
For evaluation, we used two datasets: ArtistMus (in-domain) and TrustMus (out-of-domain) (ramoneda2024role) [^3] Performance on the MusicTheoryBench (yuan2024chatmusician) and ZIQI-Eval (li2024music) benchmarks will be included in the camera-ready version. Performance on factual and contextual questions was separately measured on the ArtistMus. For TrustMus, evaluation was conducted across four categories: People, Instrument & Technology, Genre, Forms, and Theory, and Culture & History, each comprising 100 questions. All evaluations use a multiple-choice QA format, and accuracy (exact match of the predicted option A, B, C, or D) was adopted as the evaluation metric.
@@ 4.2.  Models and Experimental Setup
  • Zero-shot: GPT-4o (achiam2023gpt) (API-based), Gemini 2.5 Flash (API-based), ChatMusician (yuan2024chatmusician) (music-specific), Llama 3 (1B/3B/8B) (grattafiori2024llama), and Qwen3 (0.6B/1.7B/4B/8B) (yang2025qwen3). For Qwen3 models, we disabled the thinking mode to ensure consistent behavior.
@@
  • Rerank: We add a second-stage reranking step with the BGE reranker (bge_embedding), selecting the most relevant passages for generation.
@@
We adopt the following training configuration for both QA fine-tuning and RAG-style fine-tuning: All fine-tuned models were trained for one epoch using LoRA (hu2022lora) with 8-bit quantization. We set the following hyperparameters: batch size = 2, gradient accumulation steps = 4, learning rate = 3e-5, weight decay = 0.005, warmup ratio = 0.1, cosine learning rate scheduler (wolf-etal-2020-transformers), AdamW optimizer (loshchilov2017decoupled), LoRA rank $r=16$, scaling factor $\alpha=16$, and dropout = 0.1.
@@ 4.3.  Retriever Configurations
  • Embedding models: BM25 (robertson1994some) (sparse), Contriever (izacard2021unsupervised) (dense).
  • Reranker: bge-reranker-large (bge_embedding).
@@ 5.3.  Retriever Configuration Study
Table 7: Llama 3.1 8B Instruct (grattafiori2024llama) RAG performance on ArtistMus, by different passage size and embeddings. R indicates the use of a reranker. The best result is highlighted in bold, while the second best is underlined.
Figure 4: RAG performance and retrieval time for Wikipedia Corpus (karpukhin2020dense) and MusWikiDB.

### v2

@@ 1.  Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across diverse domains, yet their effectiveness in music-related tasks remains limited due to sparse music-specific coverage in pretraining data. This limitation is particularly evident in Music Question Answering (MQA), where users seek detailed information about artists, genres, instruments, and music history—knowledge that general-purpose LLMs often lack or hallucinate (Ramoneda et al., 2024).
Traditional approaches to domain adaptation rely on fine-tuning LLMs with domain-specific data (Jeong, 2024; Sahoo et al., 2024). However, this faces significant challenges: securing high-quality training data, escalating computational costs, and the persistent difficulty of continuously updating model knowledge as new music emerges. Retrieval Augmented Generation (RAG) (Lewis et al., 2021) offers a compelling alternative by grounding model outputs in external knowledge bases, enabling dynamic access to specialized information without retraining.
@@
Despite RAG’s success in domains such as medicine (Miao et al., 2024) and law (Wiratunga et al., 2024), its application in music remains underexplored due to the absence of music-specific retrieval resources. Existing knowledge bases like Wikipedia lack the density and structure needed for effective MQA. Similarly, current music benchmarks (Weck et al., 2024; Yuan et al., 2024; Ramoneda et al., 2024) focus primarily on music theory or multimodal understanding, leaving a critical gap in evaluating artist-centric knowledge—information central to everyday music discovery and consumption (Lee, 2010; Doh et al., 2024).
@@
Using these resources, we conduct systematic experiments showing that retrieval-augmented approaches markedly improve music QA performance. Open-source models equipped with MusWikiDB-based retrieval achieve substantial gains in factual accuracy, performing on par with or exceeding proprietary systems. Moreover, RAG-style training on (context, question, answer) triples further enhances factual recall and contextual reasoning. The framework generalizes well to out-of-domain benchmarks and yields more accurate and efficient retrieval than the general Wikipedia corpus. [^1] Code and data at https://github.com/DaeyongKwon98/ArtistMus
@@ 2.1.  Music Question Answering Benchmarks
Question Answering (QA) is a core task in Information Retrieval, involving the provision of accurate answers to given questions by retrieving relevant information from document collections (Allam and Haggag, 2012). While open-domain QA systems operate over broad knowledge bases (Rajpurkar et al., 2016; Karpukhin et al., 2020), domain-specific QA targets specialized fields such as medicine (Pal et al., 2022), law (Chalkidis et al., 2021), or music, where both documents and queries are confined to domain-specific knowledge.
Several recent benchmarks have emerged to evaluate LLM performance on music-related tasks. MuChoMusic (Weck et al., 2024) presents 1,187 audio-based multiple-choice questions assessing musical knowledge and reasoning. MusicTheoryBench (Yuan et al., 2024) contains 372 expert-validated questions evaluating advanced music theory knowledge. TrustMus (Ramoneda et al., 2024) comprises 400 questions across four domains—People, Instruments & Technology, Genres/Forms/Theory, and Culture & History—derived from The Grove Dictionary Online (Sadie and Tyrrell, 2001). ZIQI-Eval (Li et al., 2024) offers 14,000 comprehension tasks spanning 10 major topics and 56 subtopics.
However, a critical gap exists in existing benchmarks: the inadequate representation of artist-centric metadata—information crucial for everyday music listening contexts (Lee, 2010; Doh et al., 2024). Current text-only QA benchmarks lack comprehensive coverage of details that listeners frequently seek, such as complete discographies, collaboration networks, creative evolution across albums, and career achievements. This disparity between current MQA capabilities and practical information demands motivates the development of artist-focused evaluation resources.
@@
Domain adaptation of LLMs traditionally relies on fine-tuning with domain-specific data (Jeong, 2024; Sahoo et al., 2024; Satterfield et al., 2024). However, this approach faces challenges: securing high-quality training data, escalating computational costs with model scale, and difficulties in continuously updating model knowledge.
Retrieval Augmented Generation (RAG) (Lewis et al., 2021) offers an alternative by augmenting LLMs with external knowledge retrieval. Rather than relying solely on parametric memory, RAG retrieves relevant passages from external databases during inference to ground responses in factual context. Prior work demonstrates that RAG improves factual accuracy, reduces hallucination, enables transparent source verification, and provides greater scalability and cost-efficiency compared to fine-tuning alone (Borgeaud et al., 2022; Yasunaga et al., 2022; Wang et al., 2023).
While RAG has been successfully applied in general-purpose (Ram et al., 2023) and domain-specific settings such as law (Wiratunga et al., 2024) and medicine (Miao et al., 2024), its application in the music domain remains underexplored. Music QA requires both factual recall (e.g., artist names, compositions, release dates) and contextual reasoning (e.g., cultural influences, historical relations), making it an ideal testbed for evaluating RAG’s effectiveness in culturally rich, knowledge-intensive domains.
Recent work has also explored RAG-style fine-tuning, where models are trained on (context, question, answer) triples rather than traditional (question, answer) pairs (Shuster et al., 2021). This approach explicitly teaches models to integrate supporting passages, potentially improving their ability to leverage external evidence. However, systematic evaluation of this strategy within specialized cultural domains like music has been limited.
@@ 3.  Dataset
Table 3: MusWikiDB and Wikipedia Corpus (Karpukhin et al., 2020) statistics.
@@ 3.1.  MusWikiDB
To construct the RAG database from MusWikiDB_raw, the text was segmented into passages of up to 256 tokens with a 10% overlap between adjacent segments, based on the ablation study in Section 4.3, in order to preserve contextual continuity. For first-stage retrieval, we employed BM25 (Robertson and Walker, 1994), a classical yet highly effective algorithm for text relevance ranking, to build an efficient index for MusWikiDB. This enabled fast retrieval of relevant information during RAG-based inference, thereby improving both the accuracy and contextual relevance of the generated answers. The resulting MusWikiDB serves as a scalable and up-to-date knowledge base, enhancing RAG performance in MQA tasks and enabling the system to address complex, domain-specific music-related questions with greater precision and context.
Table 3 compares our proposed MusWikiDB with the Wikipedia corpus (Karpukhin et al., 2020). While MusWikiDB contains fewer pages (0.14M vs 3.2M) and has a smaller vocabulary size (7.5M vs 21.5M), it consists exclusively of music-specialized text information.
Example Questions of ArtistMus Metadata: Artist: Robyn Country: Sweden Debut Year: 1989 Genre: Pop Topic: Artistry Factual Question: "What vocal range does Robyn have?" A. Alto B. Soprano C. Tenor D. Baritone Answer: B Contextual Question: "How did Robyn’s approach to music change while writing her eighth studio album, Honey, according to the text?" A. Robyn explored music that was hypnotic and lacked a traditional structure. B. Robyn aimed to replicate the success of her previous hits. C. Robyn focused on creating more mainstream pop hits. D. Robyn returned to her R&B roots with a focus on tidy pop songs. Answer: A
@@ 3.2.  ArtistMus
While previous benchmarks have mainly addressed general music understanding, none has focused on music metadata, particularly the artist, which is crucial in music listening contexts (Lee, 2010; Doh et al., 2024). Therefore, we introduce ArtistMus, a multiple-choice benchmark dataset on music artists, to evaluate the performance of LLMs in artist-related QA using artist-specific data from MusWikiDB.
@@ 3.2.1.  Data Collection
For genre normalization (Schreiber, 2015), all genre labels were converted to lowercase, and spaces, hyphens (-), and slashes (/) were removed. We used the 48 root genres defined in (Schreiber, 2016) as our initial genre taxonomy. After filtering to retain data corresponding to the 300 most frequent genres, each was mapped to one of 20 final genre labels.
To extract artists’ regional information, we provided the first paragraph (abstract) of each Wikipedia page to Llama 3.1 8B Instruct (Grattafiori et al., 2024), prompting it to identify the country associated with each artist. The list of valid countries was obtained from the pycountry [^2] https://pypi.org/project/pycountry/ library.
@@ 3.2.2.  Data Selection
Figure 2 illustrates the regional distribution of the artists included in ArtistMus. [^3] You can see larger figure at Figure 5. Unlike many existing datasets that are heavily biased toward the U.S. and Europe, ArtistMus deliberately incorporates artists from 163 countries (or regions), ensuring broad global diversity and representation. Approximately 43% of the artists originate from outside the U.S. and Europe, indicating balanced representation beyond Western and European cultural centers.
@@
For each artist, we generated one factual and one contextual question to evaluate the LLM’s factual accuracy and contextual reasoning ability. To construct these questions, we provided GPT-4o (Achiam et al., 2023) with the corresponding section text. Factual questions target verifiable details such as dates, names, and events, whereas contextual questions require reasoning or synthesis across multiple pieces of information within the passage.
@@ 3.2.3.  Data Validation
To ensure benchmark quality, we applied a two-stage validation process. We validate the generated questions based on two criteria: Music Relevance and Faithfulness. For Music Relevance, questions that did not pertain to musical aspects were excluded except important details such as the artist’s birthplace. For Faithfulness, we prompted the LLM to verify whether the question and answer could be derived from the provided text. Finally, 1,000 multiple-choice questions passing final human validation were generated. We adjusted the placement of correct answers to ensure an even distribution, assigning 250 instances to each option.
@@ 4.1.  Benchmarks
For evaluation, we used two datasets: ArtistMus (in-domain) and TrustMus (out-of-domain) (Ramoneda et al., 2024) [^4] Performance on the MusicTheoryBench (Yuan et al., 2024) and ZIQI-Eval (Li et al., 2024) benchmarks will be included in the camera-ready version. Performance on factual and contextual questions was separately measured on the ArtistMus. For TrustMus, evaluation was conducted across four categories: People, Instrument & Technology, Genre, Forms, and Theory, and Culture & History, each comprising 100 questions. All evaluations use a multiple-choice QA format, and accuracy (exact match of the predicted option A, B, C, or D) was adopted as the evaluation metric.
@@ 4.2.  Models and Experimental Setup
  • Zero-shot: GPT-4o (Achiam et al., 2023) (API-based), Gemini 2.5 Flash (API-based), ChatMusician (Yuan et al., 2024) (music-specific), Llama 3 (1B/3B/8B) (Grattafiori et al., 2024), and Qwen3 (0.6B/1.7B/4B/8B) (Yang et al., 2025). For Qwen3 models, we disabled the thinking mode to ensure consistent behavior.
@@
  • Rerank: We add a second-stage reranking step with the BGE reranker (Xiao et al., 2023), selecting the most relevant passages for generation.
@@
We adopt the following training configuration for both QA fine-tuning and RAG-style fine-tuning: All fine-tuned models were trained for one epoch using LoRA (Hu et al., 2022) with 8-bit quantization. We set the following hyperparameters: batch size = 2, gradient accumulation steps = 4, learning rate = 3e-5, weight decay = 0.005, warmup ratio = 0.1, cosine learning rate scheduler (Wolf et al., 2020), AdamW optimizer (Loshchilov and Hutter, 2017), LoRA rank $r=16$, scaling factor $\alpha=16$, and dropout = 0.1.
@@ 4.3.  Retriever Configurations
  • Embedding models: BM25 (Robertson and Walker, 1994) (sparse), Contriever (Izacard et al., 2021) (dense).
  • Reranker: bge-reranker-large (Xiao et al., 2023).
@@ 5.3.  Retriever Configuration Study
Table 7: Llama 3.1 8B Instruct (Grattafiori et al., 2024) RAG performance on ArtistMus, by different passage size and embeddings. R indicates the use of a reranker. The best result is highlighted in bold, while the second best is underlined.
Figure 4: RAG performance and retrieval time for Wikipedia Corpus (Karpukhin et al., 2020) and MusWikiDB.